OpenStudy (anonymous):

find an exponential function of the form y=ab exponent x whose graph passes throughthe points (2,1) and (3,2)

ganeshie8 (ganeshie8):

lets call the function \(y = ab^x\) it goes thru (2,1) => \(1 = ab^2\) ---------------(1) it goes thru (3,2) => \(2 = ab^3\) ---------------(2) (2)/(1) => b=2 now substitute b value in (1) and find the value of a

ganeshie8 (ganeshie8):

can u try the rest of it @bri4life14

OpenStudy (anonymous):

Before we do anything, we must confirm what he means by ab to the power of x. Do you mean:\[y=(ab)^x\] OR \[y=ab^x\] @bri4life14

ganeshie8 (ganeshie8):

depends.. if u are off from academic u will be all uncertain. but the general stuff we see everyday - we dont need confirmation as we see these questions everyday

OpenStudy (anonymous):

im sorry i dont know what im doing :(

OpenStudy (goformit100):

@erdog82

OpenStudy (anonymous):

\[y=ab^x\] you need \(a\) and \(b\) we can find \(b\) first \[y=ab^x\]and since \((2,1)\) is on the graph you know \[1=ab^2\]

OpenStudy (anonymous):

since \((3,2)\) is on the graph you also know \[2=ab^3\] now you can find \(b\) by division, since \[\frac{ab^3}{ab^2}=b\] you know \[\frac{2}{1}=2=b\]

OpenStudy (anonymous):

so now you have \[y=a\times 2^x\] and now you can find \(a\) the same way as above